24/6 = n/5 what does n equal? Please work out the problem.
Answer by richwmiller(17219)

Exactly you should work out the problem.
Multiply both sides first by 5 then by 6
(24/6)*6=6n/5
now by 5
5*24*6/6=5*6n/5
now look for numbers to cancel out such as 6/6 and 5/5
and then multiply out the remaining numbers